### v1

# A Quantitative Evaluation of the Expressivity of BMI, Pose and Gender in Body Embeddings for Recognition and Identification

Abstract
[…]
I INTRODUCTION
Deep learning models are primarily designed to learn specific target attributes during training; however, they frequently encode unintended, image-related attributes that can introduce significant biases into the system. In the domain of biometrics, particularly face recognition, Hill et al. [18] demonstrated that deep networks form identity representations that inherently cluster based on gender. Moreover, these identity embeddings have been shown to encode other latent characteristics such as pose, age, and lighting conditions [18, 29, 34]. The presence of these latent attributes can significantly influence the accuracy of recognition algorithms, as they may inadvertently affect model predictions [15, 26]. Evaluating these biases or understanding the correlation between attributes and network features require a systematic analysis of how these attributes are embedded and how they influence model behavior. A deeper understanding of these phenomena necessitates investigating how facial or body attributes are encoded in identity representations and how they shape predictive outcomes. In this context, Dhar et al. [12] introduced the concept of expressivity, for face recognition, a metric that quantifies the relationship between learned network features and specific attributes, thereby enhancing the interpretability of face recognition models. Building on this concept, we extend the framework of expressivity to the domain of person re-identification (ReID) with the goal of evaluating how body-related features are embedded within ReID models trained primarily for identity recognition.
Person re-identification (ReID) is a well-established research area with a range of real-world applications, including smart city infrastructure for public safety and traffic management [2, 22] and autonomous driving systems for pedestrian detection and tracking [4, 41]. The primary objective of ReID is to accurately match and retrieve pedestrian identities across non-overlapping camera views, varying time frames, and distinct locations, all while addressing challenges such as pose variations, appearance diversity, and environmental conditions [16, 17, 46]. Significant progress has been achieved in improving re-identification accuracy through the development of deep learning methods, which can broadly be categorized into image-based and video-based approaches. Image-based ReID methods focus on selecting the most distinctive frame and extracting fine-grained spatial features, while video-based approaches aggregate temporal information across multiple frames to produce more robust identity representations. Recent advancements have increasingly combined these approaches, leveraging the strengths of both image-level detail and temporal consistency to achieve state-of-the-art performance. Despite these advancements, most deep learning-based ReID systems are trained to identify individuals based on visual body features, without explicitly learning specific body-related attributes. These models generate identity representations derived from body cues; however, similar to face recognition systems, ReID networks often unintentionally encode additional attributes related to body characteristics. To address this, our work conducts a comprehensive analysis of the attributes correlated with feature embeddings generated by the state-of-the-art SemReID model. As ReID systems are increasingly deployed in real-world applications, there is a growing demand for explainable and transparent models. Understanding how various attributes are encoded across internal network layers is crucial for interpreting identity predictions and identifying potential sources of algorithmic bias. The following are the conceptual and experimental contributions of our paper:
  • We present the first investigation into the encoding of body attributes within the layers of a large-scale Vision Transformer (ViT)-based foundation model for person ReID. To enhance the interpretability of large-scale deployable ReID systems, we propose a novel post-hoc framework that explains how internal representations influence identity predictions. This achievement underscores the robustness of our approach, despite the inherent complexity of the model and the diversity of the dataset.
  • In the final attention layer of the SemReID network, we observe the following order of expressivity for body attributes: BMI $>$ Pitch $>$ Yaw $>$ Gender. This ranking highlights the varying degrees of influence that different attributes have on the network’s predictions.
  • To provide a more comprehensive understanding, we analyze how feature-attribute correlations evolve across different layers and throughout the training process. This layer-wise and temporal analysis offers deeper insights into the embedding of body attributes and their impact on ReID performance.
II RELATED WORKS
Person re-identification (ReID) is a critical task in computer vision that focuses on matching individuals across non-overlapping camera views, often under varying conditions such as illumination, clothing, pose, and occlusion [16, 17, 21, 46]. Numerous works have been proposed to address this challenge in various ReID domains, such as Clothes-Changing ReID (CC-ReID)[16], video ReID[5, 19, 43, 45, 42], unconstrained ReID[11, 27, 31, 30, 47], and short-term ReID[9, 45, 40, 48]. Among these methods, the SemReID [20] approach achieves state-of-the-art (SoTA) performance across all four domains. Though the interpretability of these specific models have not been extensively explored in prior literature, several works have examined the explainability of recognition algorithms in general, primarily focusing on face recognition.
The study of bias and interpretability in biometrics has long been a subject of significant interest [36, 13, 38, 32, 33]. Schumann et al. [35] introduced an auxiliary network to enhance the performance of ReID systems by incorporating complementary information into CNN features. More recently, Myers et al. [28] demonstrated that linguistic and non-linguistic representations of body shape can provide complementary identity information, improving identification in specific applications. These methods rely on the change in predictions with respect to a concept or attribute to interpret a network’s sensitivity to the attribute. Yin et al. [44] introduced a spatial activation diversity loss to maintain interpretability in face recognition networks during training. Similarly, Kim et al. [23] proposed a generative approach utilizing representative exemplars (prototypes) to enhance model interpretability. However, as noted in [24], these methods are limited to models trained from scratch and cannot be applied to pre-trained networks or those already deployed in practice.
Post-hoc interpretability methods have also been explored to understand trained models, with TCAV (Testing with Concept Activation Vectors) [24] being one of the most influential approaches in this domain. TCAV interprets a model’s behavior by analyzing its sensitivity to user-defined concepts and achieves this by learning Concept Activation Vectors (CAVs) through training a linear classifier to distinguish activations produced by concept examples. While TCAV is effective for discrete physical concepts like color or texture, it is not suitable for assessing sensitivity to continuous or abstract attributes, such as pose angle or BMI. This limitation arises because generating negative examples (images without the concept) is challenging for omnipresent and continuous attributes. Furthermore, TCAV requires test images to belong to one of the training classes, as its computations rely on changes in logits for specific classes, rendering it incompatible with scenarios involving unseen subjects or faces. Other post-hoc methods, such as [1], utilize linear classifiers to analyze intermediate network layers, while Koh and Liang [25] proposed influence functions to measure model sensitivity to infinitesimal perturbations in training data. However, these methods are unsuitable for evaluating sensitivity to physical attributes like pose or orientation, as they primarily focus on localized perturbations. Saliency- and attention-based approaches [37, 8] offer another class of interpretability methods, generating attention maps to visualize spatial regions influencing model predictions. For person ReID, Chen et al. [10] introduced a learnable, pluggable interpreter for CNN-based models, which decomposes image-pair distances into attribute-based contributions and visualizes attention maps for discriminative attributes. While effective, this method relies on attribute-guided metric distillation, involves computationally intensive post-hoc operations, and is not adapted for transformer-based backbones. Moreover, while saliency maps can highlight spatial regions, they cannot assess sensitivity to abstract or non-localized concepts, such as BMI or pitch/yaw. Few studies explicitly address the interpretability of face recognition models on the basis of attribute hierarchy, for example Hill et al. [18] demonstrated a hierarchy within facial feature representations, showing that identity is nested under sex, illumination is nested under identity, and viewpoint is nested under illumination. Dhar et al. [12] further extended this analysis by conducting an expressivity-based evaluation of facial attributes, identifying a hierarchy where age has the greatest influence, followed by sex and yaw.
We propose the use of expressivity as a measure for person ReID systems, quantifying the predictability of an attribute within a given set of features extracted using the model. Unlike prior methods, expressivity can be computed for both categorical and continuous attributes, allowing for a direct comparison of attribute predictability. Additionally, expressivity can be calculated agnostic to the ReID network backbone. In this work, we demonstrate results using a SoTA Vision Transformer (ViT)-based ReID model. This framework provides valuable insights into how body-related features are embedded within ReID models and their influence on performance, laying the groundwork for more interpretable and explainable ReID systems.
III Proposed Method
Our approach attempts to find the correlations between the learnt features by a state-of-the-art (SoTA) body recognition model and sensitive attributes, as highlighted in Figures 1 and 2. The predictability of attributes from a given set of body descriptors reflects the amount of attribute-relevant information encoded within those descriptors. To quantify this information, we employ Mutual Information (MI) as shown in Equation 1. MI is a fundamental quantity for measuring the relationship between random variables, indicating how much knowledge of one variable reduces uncertainty about the other. By estimating the MI between features learned by the body recognition model and their corresponding sensitive attributes, we assess the degree to which these descriptors encode attribute information. Since MI captures non-linear statistical dependencies between variables and is applicable to both categorical and continuous attributes, this approach provides a unified and consistent measure across attribute types. To develop a general-purpose estimator, we utilize the widely recognized formulation of MI as the Kullback-Leibler (KL) divergence (Kullback, 1997) between the joint distribution and the product of the marginal distributions of two random variables $X$ and $Z$, as expressed in Equation 2.
$I(X;Z)=\int_{\mathcal{X}\times\mathcal{Z}}\log\frac{d\mathbb{P}_{XZ}}{d\mathbb{P}_{X}\otimes\mathbb{P}_{Z}}d\mathbb{P}_{XZ}$ (1)
$I(X;Z)=D_{KL}\left(\mathbb{P}_{XZ}\|\mathbb{P}_{X}\otimes\mathbb{P}_{Z}\right)$ (2)
III-A Problem Setup
Figure 1: Integrating the MINE block in the ViT based SemReID [20] backbone to compute the expressivity of features with respect to attributes such as BMI, gender, pitch and yaw. The internal structure of the MINE block is shown in the next figure.
Figure 2: The internal structure of the MINE block employs a simple MLP with two hidden layers to compute the expressivity of $m$-dimensional features $F$. By augmenting these features with an attribute vector $A$, the input to the network is extended to $(m+1)$-dimensions.
Our dataset comprises body images of different individuals captured under varying conditions and at different distances. Each image is annotated with an identity label and several sensitive attributes, including gender ($g$), height ($h$), weight ($w$), body mass index ($BMI$), which is computed from $h$ and $w$, as well as pitch angles ($p$) and yaw angles ($y$). These attributes collectively form a diverse set of information, enabling a comprehensive analysis of how sensitive attributes are encoded in the learned features. We denote the set of learned feature descriptors as $\mathbf{F}$, and the corresponding sensitive attributes as $A$. The primary objective of this analysis is to quantify and explore the correlations between $\mathbf{F}$ and $A$ using Mutual Information Neural Estimation (MINE). Specifically, we aim to estimate the MI, denoted as $I_{\theta}(\mathbf{F},\mathbf{A})$, to gain insights into how effectively the learned features capture attribute-relevant information.
To achieve this, each image $x_{i}$ undergoes a series of preprocessing steps before being passed through the SemReID model. The model extracts feature descriptors $f_{i}$, where $f_{i}\in\mathbb{R}^{m}$, representing the encoded identity and attribute information for each image. These descriptors are then concatenated to form the feature matrix $\mathbf{F}=[f_{1},f_{2},\dots,f_{n}]^{T}$, where $F\in\mathbb{R}^{n\times m}$. The sensitive attribute vector $A\in\mathbb{R}^{n\times 1}$, containing information such as gender, pose, and identity, is then combined with $F$ to form an augmented matrix $\mathbf{X}=[\mathbf{F}|\mathbf{A}]$. This augmented matrix $X$ is subsequently used by the MINE network to estimate the MI between $\mathbf{F}$ and $\mathbf{A}$.
MINE employs a neural network-based approach to approximate the MI, enabling us to compute $I_{\theta}(\mathbf{F},\mathbf{A})$ effectively, even in high-dimensional feature spaces. By leveraging this approach, we can evaluate the extent to which the learned feature descriptors $\mathbf{F}$ encode information relevant to the sensitive attributes $\mathbf{A}$. This analysis provides valuable insights into the relationship between the network’s internal representations and sensitive attributes, helping to understand potential biases and attribute-specific influences in the model’s learned features.
III-B Attributes and Their Relevance
We compute the expressivity of four annotated attributes: $g$, $BMI$, $p$ and $y$ in the extracted features. In Figure 3, we verify that the dataset we utilize shows enough variation with respect to these attributes, so that we can ensure that expressivity (which is a lower bound estimate of MI) is an accurate model for the corresponding attributes. When considering g, the vector $\mathbf{A}$ is a discrete vector having a value of 1 if the gender is male and 0 if female while for $BMI$, y and p (in degrees) the values of the vector values are continuous. These attributes play a vital role in person ReID tasks, as they influence the expressivity of the model’s learned features. Attributes like $BMI$ or pose are particularly challenging to disentangle from identity for this task, making them ideal for analyzing feature relevance.
Figure 3: Attribute distribution and counts in the BRIAR dataset indicate sufficient variation across the attributes of interest.
III-C Expressivity of Body Features
Understanding the expressivity of learned features in deep networks is critical for tasks that rely on nuanced feature representations, such as person re-identification (ReID). Tishby and Zaslavsky [39] introduced the concept of utilizing MI as a quantitative metric to assess how well information is retained or transformed across the layers of a deep network. MI measures the dependency between random variables, offering insights into the trade-offs between compression and informativeness at various stages of a network. By quantifying MI, one can directly evaluate how effectively the network balances these competing objectives.
However, estimating MI for high-dimensional continuous variables is computationally challenging due to the need to compute probability density functions of the underlying distributions. Traditional methods often rely on discretization or kernel density estimation, both of which suffer from scalability issues as dimensionality increases. To overcome this, Belghazi et al. [3] proposed Mutual Information Neural Estimation (MINE), a scalable framework that approximates MI using a neural network. This approach bypasses the need for explicit density computations by optimizing a neural network-based lower bound of MI, making it suitable for high-dimensional and complex datasets.
The MI between learned feature descriptors $\mathbf{F}$ and sensitive attributes $\mathbf{A}$ is a crucial metric in evaluating the expressivity of the learned features. In the context of this work, $\mathbf{F}$ represents the feature embeddings produced by the SemReID model, while $\mathbf{A}$ denotes associated sensitive attributes such as gender, pose, and identity. The MI is mathematically defined as:
$I_{\theta}(\mathbf{F},\mathbf{A})=\sup_{\theta\in\Theta}\mathbb{E}_{P_{FA}}\left[T_{\theta}(f,a)\right]-\log\mathbb{E}_{P_{F}\otimes P_{A}}\left[e^{T_{\theta}(f)}\right],$ (3)
where $T_{\theta}(f,a)$ is a neural network parameterized by $\theta$, designed to approximate the MI. The joint expectation $\mathbb{E}_{P_{FA}}\left[T_{\theta}(f,a)\right]$ measures the network’s output when conditioned on the true joint distribution of features and attributes. In contrast, the term $\log\mathbb{E}_{P_{F}\otimes P_{A}}\left[e^{T_{\theta}(f)}\right]$ normalizes the MI estimate to ensure it captures only the dependency between $\mathbf{F}$ and $\mathbf{A}$, excluding any bias from their marginal distributions.
Computational Steps for Mutual Information Estimation
Step 1: Joint Expectation Approximation.
The first term, $\mathbb{E}_{P_{FA}}\left[T_{\theta}(f,a)\right]$, quantifies the degree of dependency between $\mathbf{F}$ and $\mathbf{A}$ by evaluating the network output over their joint distribution. In practice, this expectation is approximated over minibatches of data as:
$\mathbb{E}_{P_{FA}}\left[T_{\theta}(f,a)\right]\approx\frac{1}{b}\sum_{i=1}^{b}T_{\theta}(f_{i},a_{i}),$ (4)
where $b$ is the batch size, and $f_{i}$ and $a_{i}$ are the $i$-th feature vector and attribute value in the batch, respectively. This term essentially aggregates the network’s outputs for each feature-attribute pair, capturing their joint statistics.
Step 2: Marginal Expectation Approximation.
The second term, $\log\mathbb{E}_{P_{F}\otimes P_{A}}\left[e^{T_{\theta}(f)}\right]$, ensures that the MI estimate reflects only the mutual dependency, independent of marginal distributions. It is computed by approximating the expectation of the exponential of the network’s output under the product of marginals:
$\mathbb{E}_{P_{F}\otimes P_{A}}\left[e^{T_{\theta}(f)}\right]\approx\frac{1}{b}\sum_{i=1}^{b}e^{T_{\theta}(f_{i})}.$ (5)
This term prevents the MI estimate from over-representing trivial correlations caused by the underlying marginal distributions.
Step 3: Objective Function Formulation.
The MI lower bound is approximated as the difference between the joint and marginal expectations, yielding the following objective function:
$V(\theta)=\frac{1}{b}\sum_{i=1}^{b}T_{\theta}(f_{i},a_{i})-\log\left(\frac{1}{b}\sum_{i=1}^{b}e^{T_{\theta}(f_{i})}\right).$ (6)
Maximizing $V(\theta)$ corresponds to maximizing the MI lower bound, thus enabling the neural network to learn representations that effectively capture the mutual dependency between features and sensitive attributes.
Step 4: Loss Function and Optimization.
To train the neural network $T_{\theta}$, the negative of the objective function is used as the loss:
$L(\theta)=-V(\theta).$ (7)
The gradient of this loss function with respect to the parameters $\theta$ is computed as:
$\nabla_{\theta}L(\theta)=-\left(\mathbb{E}_{P_{FA}}\left[\nabla_{\theta}T_{\theta}\right]-\mathbb{E}_{P_{F}\otimes P_{A}}\left[\nabla_{\theta}e^{T_{\theta}}\right]\right).$ (8)
This gradient is then used to iteratively update the network parameters using gradient descent. To mitigate biases introduced by minibatch sampling, an exponential moving average of the gradients is applied during optimization.
In the context of person ReID, this framework is particularly valuable for understanding the expressivity of the feature descriptors generated by the model. The neural network $T_{\theta}$ is trained to approximate the MI between the learned features $\mathbf{F}$ and sensitive attributes $\mathbf{A}$. By iteratively computing joint and marginal expectations and updating $\theta$, the MI provides a robust metric to quantify how much attribute-relevant information is encoded in the features. At convergence, the MI value reflects the extent to which the model’s representations capture sensitive attribute information, offering insights into the expressivity and fairness of the learned features.
Algorithm 1 Expressivity Computation on learnt representations
1:Layer $L$, set of $n$ images $I$, attribute vector $\mathbf{A}\in\mathbb{R}^{n\times 1}$
2:Expressivity measure
3:Initialize $E\leftarrow[]$triangleright To store expressivity values
4:Extract features $\mathbf{F}\leftarrow[f_{1},f_{2},\dots,f_{n}]^{T}$ from $L$ after a particular epoch for all $i\in I$
5:Concatenate the features and attributes: $\mathbf{X}\leftarrow[\mathbf{F}|\mathbf{A}]$triangleright Augmentation step
6:for $\text{iteration}=1$ to $M$ do
7:     Initialize MINE network $T_{\theta}$ based on the dimensions of $\mathbf{X}$
8:     Compute expressivity score: $e\leftarrow\text{MINE}(\mathbf{X})$
9:     Append score: $E\leftarrow E\cup\{e\}$
10:end for
11:return $\text{Expressivity}\leftarrow\text{Average}(E)$
IV EXPERIMENTS
IV-A Dataset and Settings
Figure 4: Attribute annotated exemplar images from the BRIAR dataset.
We use the BRIAR 1-5 dataset [11] which has 382,229 images and 170,522 videos of which we extract 704,999 frames of 2,077 unique identities for our work. In our curated data subset, there are 887 males and 1,190 unique female subjects. BRIAR is a massive unconstrained person re-identification dataset containing over 40,000 videos and 1 million images captured across challenging real-world conditions, including varying clothes, distances (100m, 200m, 400m, 500m, 600m, 800m, 1km), altitudes (close range to UAV), and environmental factors like occlusion, blur, and physical turbulence. The dataset consists of 5 variants (BRIAR-1, 2, 3, 4, and 5), each with increasing complexity in terms of number of identities, distractors, and capturing conditions, making it a comprehensive benchmark for testing re-identification systems in unconstrained scenarios. Some example images from the dataset along with the required attribute annotation is shown in Figure 4.
IV-B Integration of MINE with SemReID
SemReID is a self-supervised person re-identification model that introduces a novel Local Semantic Extraction (LSE) module. This module leverages keypoint predictions to guide an interactive segmentation model, Segment Anything Model (SAM), in extracting precise local semantic masks for various body parts. These masks enable the model to isolate and extract fine-grained biometric features that enhance its ability to discern unique identities. Trained using a teacher-student framework, the SemReID model incorporates multiple loss functions to ensure the learned representations are robust and transferable. This approach simplifies the inference process by utilizing only the teacher network, paired with a single linear layer, to perform efficient re-identification. During inference, the model processes input data—including both still images and video frames—through the teacher encoder. By concatenating global and local features, the SemReID model effectively generalizes across diverse re-identification domains, negating the need for domain-specific adaptations or fine-tuning.
To integrate MINE into the SemReID pipeline, an auxiliary neural network is employed to estimate MI by maximizing the Donsker-Varadhan (DV) lower bound. As described earlier, MINE provides a robust framework for estimating MI between learned feature representations and sensitive attributes, such as gender, pose, or identity. The implementation of this integration is detailed in Algorithm 1, which outlines the process for computing the expressivity of the learned representations.
The process begins with a dataset consisting of $n$ images, each associated with a corresponding sensitive attribute vector $\mathbf{A}\in\mathbb{R}^{n\times 1}$. Features for these images are extracted from a specific layer $L$ of the SemReID model after a training epoch. These features, denoted as $\mathbf{F}=[f_{1},f_{2},\dots,f_{n}]^{T}$, represent both global and local information captured by the model. To prepare the input for the MINE network, the features and attributes are concatenated to form an augmented matrix $\mathbf{X}=[\mathbf{F}|\mathbf{A}]$. The MINE network $T_{\theta}$ is then initialized based on the dimensions of $X$. The network architecture, described in Figure 2, is a multi-layer perceptron (MLP) comprising two hidden layers with 256 and 128 units, respectively. These layers use Exponential Linear Unit (ELU) activations, chosen for their ability to ensure stable training and mitigate issues such as vanishing gradients. The MLP computes $T_{\theta}$, which is used to estimate the MI by maximizing the DV lower bound. To compute the expressivity of the learned representations, the process iterates for $M$ iterations. In each iteration, the MINE network processes the augmented matrix $\mathbf{X}$ to compute an expressivity score $e$. This score reflects the MI between the feature representations and the sensitive attributes. The scores from all iterations are stored in a list $E$, and the final expressivity measure is obtained by averaging these scores, as described in Algorithm 1.
This integration of MINE with SemReID provides a principled approach to evaluating the expressivity of learned representations. By quantifying the MI between features and sensitive attributes, it becomes possible to assess the extent to which the model captures attribute-relevant information. Furthermore, the iterative nature of the process ensures robust estimates of expressivity while minimizing biases introduced by stochastic factors during training. Importantly, this integration is computationally efficient, as the MINE network operates on features extracted from the SemReID model, making it scalable.
IV-C Hierarchical and Temporal Analysis of Attribute Influence
To achieve a comprehensive end-to-end analysis of attribute influence within our framework, we evaluate feature and attribute correlations both hierarchically across the model’s layers and temporally throughout the training process. The SemReID model employs a Vision Transformer (ViT) backbone with 12 attention layers, offering a powerful architecture for capturing both global and local semantic information.
For the hierarchical analysis, we extract features from multiple attention layers, specifically layers 2, 4, 6, 9, and 12. These layers are chosen to provide a granular understanding of how attribute correlations evolve at different depths of the network. By analyzing the intermediate layers, we gain insights into the progression of feature learning, starting from the lower layers that capture basic patterns and spatial information to the deeper layers where high-level semantic representations emerge. This hierarchical evaluation allows us to observe the model’s ability to encode sensitive attributes at various stages of the feature extraction pipeline and how these representations are refined.
For the temporal analysis, we assess the evolution of feature-attribute correlations across training epochs. Building on prior work that established 11 epochs as the optimal training duration for identity recognition, we extend this analysis to include intermediate epochs. Specifically, we extract features from layers 2, 4, 6, 9, and 12 at training epochs 1, 3, 5, 8, and 11. This temporal sampling strategy allows us to investigate the dynamic changes in feature-attribute correlations as the model progresses through training. By observing earlier epochs, such as epochs 1 and 3, we capture the initial stages of representation learning, where the model begins to encode attribute information. In contrast, later epochs, such as 8 and 11, reveal how the model converges and solidifies its understanding of attributes in its learned representations.
This dual analysis framework examining both the hierarchical layers and temporal progression offers a detailed perspective on how attribute information is processed and encoded within the SemReID model. Hierarchical analysis highlights the architectural contribution of different layers to feature representation, while temporal analysis sheds light on the dynamic learning trajectory of attribute representations over time. Together, these analyses enable us to uncover trends and patterns in attribute encoding, providing valuable insights into the model’s learning process in mitigating or amplifying attribute correlations.
IV-D Implementation Details
We initialize the MINE approximation network based on the input dimensions of the augmented matrix. As outlined earlier, the network is trained to estimate the lower bound of the MI between features $F$ and attribute $A$. For this purpose, we employ a simple multi-layer perceptron (MLP) architecture, illustrated in Figure 2, to compute $T_{\theta}$. The MLP consists of two hidden layers with 256 and 128 units, respectively, followed by ELU activation functions. This architecture is used consistently across all experiments. The MINE network is trained using the Adam optimizer with a learning rate of 0.001 and a batch size of 100. When different sets of features are used for $F$, only the input layer’s dimensions are adjusted to match the feature dimensionality. The network is trained until the loss in Equation (7) converges, and the expressivity is calculated as described in Algorithm 1. For all experiments, we set $M=16$. SemReID uses Vision Transformer (ViT) [14] variants as the backbone. Images are processed at $384\times 128$ resolution with a single forward pass. The inference pipeline uses a dual-stream semantic parsing strategy, where global features are processed at $768$ dimension and local semantic features at $3\times 768$ dimension where $3$ is the number of local areas. In our features they are face, upper body, and lower body. The final local features is the average of all local areas. We implement multi-crop feature augmentation [6, 7] using a cascade of $M=2$ global views and $N=3$ local views, followed by $L=12$ cross-attention layers for feature alignment. The identity embedding is computed through a BN layer, optimized for minimal latency while maintaining discriminative capabilities. We extract the features from this trained network which have a dimension of 1536 and pass it through the MINE block after concatenating it with the attribute vector to estimate MI.
V RESULTS AND DISCUSSIONS
We present our observations on the correlation between features and attributes in the SemReID model using MINE. The discussion is structured into three subsections: the first examines feature-attribute correlations within the hierarchical feedforward pass, while the second analyzes the evolution of these correlations throughout the training process and the third explains the advantages of our method.
TABLE I: Summary of values for Gender, BMI, Yaw, and Pitch across different layers.
Layer Gender BMI Yaw Pitch
Attention Layer 2 $6.57\times 10^{-3}$ 0.005 0.13 0.42
Attention Layer 4 0.003 0.149 0.283 0.523
Attention Layer 6 0.004 0.72 0.73 0.73
Attention Layer 9 0.005 0.77 0.73 0.73
Attention Layer 12 0.03 1 0.10 0.41
Figure 5: Expressivity trend of gender, yaw, pitch and BMI in input image over layer-wise learnt features from SemReID.
V-A Layer-Wise Progression in Feedforward Processing
It is crucial to emphasize that the model was trained exclusively on identity labels, without explicit supervision for BMI, gender, or pose attributes. Given that our architecture employs a ViTbase backbone, it consists of 12 transformer encoder blocks, each containing an attention layer. To analyze the evolution of feature representations, we compute the MINE at various depths—specifically after layers 2, 4, 6, 9, and 12—as summarized in Table 1 and depicted in Figure 5. The key observations are as follows:
  • In the initial layers (e.g., Layer 2), MI between the learned representations and pitch/yaw attributes is relatively high, whereas gender and BMI exhibit negligible expressivity. This suggests that lower layers predominantly encode coarse spatial and geometric features, which are crucial for pose estimation.
  • As we progress deeper into the ViT_base network, BMI expressivity increases rapidly, suggesting that this is a fine grained feature and is important for the model. By Layer 6, all attributes exhibit substantial MI with the learned features. However, gender expressivity remains comparatively lower, implying that it is not significantly encoded in the model’s learned representations.
[…]
Thus, for the semReID model trained purely on identity labels, BMI emerges as the most correlated attribute, followed by pose, with gender being the least correlated. This conclusion aligns with intuitive human perception, as body recognition inherently relies on a person’s shape and pose as primary cues. From a perceptual standpoint, humans often associate body identity with physical attributes such as body shape, proportions, and posture, making BMI and pose naturally dominant in recognition tasks.
We want to clarify at this stage that, the data processing inequality (DPI) [23] states that for three random variables $P$, $Q$, and $R$ forming a Markov chain $P\rightarrow Q\rightarrow R$, the MI satisfies $MI(P,Q)\geq MI(P,R)$. This principle formalizes the idea that no processing of data can increase MI. In the context of neural networks, let $P$ represent a random variable (e.g., BMI, pose, or identity), and $Q$ and $R$ represent features extracted at different layers of the network, where $R$ is a deterministic function of $Q$ (i.e., $R$ corresponds to a deeper layer). Since $R$ is derived from $Q$, $P$, $Q$, and $R$ form a Markov chain, and the DPI implies that the MI between $P$ and the features cannot increase as we go deeper into the network. However, the expressivity results are not strictly monotonically decreasing, which might seem to contradict the DPI. This apparent inconsistency arises because, in our work, expressivity refers to the alignment between feature representations and a given attribute, rather than MI in the strict sense of information theory. As noted in [14], the feature representations in this context are more closely tied to the predictability of an attribute rather than its information-theoretic content. Therefore, expressivity in this work reflects the strength of the relationship between features and attributes, rather than their theoretical MI. This distinction allows us to interpret the expressivity results as a measure of the model’s ability to capture attribute-specific representations, which can vary across layers.
Figure 6: Expressivity trend of gender, yaw, pitch and BMI in input image over epoch-wise learnt features from SemReID.
TABLE II: Summary of values for Gender, BMI, Yaw, and Pitch across different epochs.
Epoch Gender BMI Yaw Pitch
Epoch 1 0.012 0.87 1.3 0.84
Epoch 3 0.402 0.305 0.009 0.004
Epoch 5 0.004 0.76 $6\times 10^{-3}$ $4\times 10^{-3}$
Epoch 8 0.012 0.80 0.05 0.22
Epoch 11 0.03 1 0.10 0.41
V-B Progression of Training
To understand how attribute-feature correlations evolve over training, we analyze the expressivity of gender, yaw, pitch, and BMI in the final attention layer across different training epochs. This is summarized in Table II and visualized in Figure 6. At each epoch, the feature representations are 1536-dimensional, and attribute-wise expressivity is computed using Algorithm 1. We observe that:
[…]
  • A significant reduction in yaw expressivity is observed, while BMI expressivity remains relatively stable. This suggests that yaw-related variations are gradually suppressed as the model prioritizes more identity-relevant features. Notably, both yaw and gender exhibit minimal correlations with the learned representations, implying that the model is learning to be invariant to these attributes in its deeper layers. Conversely, BMI remains a highly correlated feature, followed by pitch, reinforcing its role as a more identity-relevant attribute.
[…]
These findings indicate that the training process makes feature representations increasingly invariant to certain body attributes, such as yaw and gender, while preserving information related to BMI and to some extent pitch.
V-C Advantages of Expressivity for Person ReID
In this subsection we justify the usage of MINE over other existing methods. The key reasons can be summarized as:
  1. 1.
    Enables Attribute Comparison Across Features: In the context of person ReID, comparing the significance of various attributes within a network’s feature set is critical for understanding how identity information is organized. Expressivity facilitates this by normalizing the error rates of attributes to a common scale, as illustrated in Figures 5 and 6. This standardization enables direct comparison of attribute contributions, unlike conventional metrics influenced by varying scales.
  2. 2.
    Supports Both Discrete and Continuous Attributes: This work focuses on body attributes for ReID, expressivity is versatile and can be applied to both discrete (e.g., gender) and continuous (e.g., pitch angle) attributes. For instance, the expressivity of a concept such as gender can be computed if a binary attribute vector $A$ (indicating 1 for male and 0 for female or vice versa) is available. Unlike TCAV [24], which is well-suited for discrete attributes, expressivity extends to continuous concepts like pose angle or BMI, where identifying negative examples (images where the concept is absent) is challenging. This capability is particularly valuable for person ReID, where continuous attributes often play a critical role.
  3. 3.
    Independent of Training Identities: Previous methods require computing changes in logits, which limits their applicability to images belonging to training identities. In contrast, expressivity does not rely on logit changes or training identity classes. This independence makes it an effective tool for analyzing unseen attributes not explicitly included during training.
VI Conclusion
We propose a method to quantify the information a ViT-based person ReID network learns about various attributes without being explicitly trained on them, by analyzing their expressivity on learnt features. This enables us to identify attributes most relevant to identity recognition across hierarchical layers and training epochs. Several important findings emerge from our investigation: (1) BMI consistently shows the highest expressivity, especially in deeper layers (e.g., layer 12) and later training stages (e.g., epoch 11), making it the most critical attribute for identity recognition even without explicit labels. (2) Attributes like yaw and pitch are expressive in mid-layers (e.g., layers 4 and 6) but lose influence in deeper layers. (3) Temporally, BMI expressivity increases throughout training, while yaw and pitch decline sharply, with yaw showing the steepest drop. Gender, notably, has minimal correlation with learned features. These findings highlight BMI as the most significant attribute, followed by yaw and pitch and gender for the person ReID task. However, since expressivity approximates MI, it is influenced by entropy and attribute label distribution, potentially affecting cross-attribute comparisons which is an inherent limitation of all MI-based approaches.
VII ACKNOWLEDGMENTS
SH and RC are supported by the BRIAR project. This research is based upon work supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via [2022-21102100005]. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U. S. Government. The US. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
VIII Ethical Impact Statement
This research utilizes the BRIAR dataset, which was collected by the Test and Evaluation team under strict consent guidelines and with approval from the appropriate ethical review board. All subjects involved provided informed consent for their participation, including the use of their images in research publications and figures. Our study aims to analyze how demographic and physiological covariates such as body mass index and gender may introduce biases in body-based recognition systems. By systematically examining these potential biases, we seek to contribute to the broader discourse on fairness, accountability, and transparency in biometric technologies. Understanding and mitigating bias in recognition systems is critical for ensuring equitable outcomes across diverse populations. This research aligns with the outlined ethical principles by advocating for responsible AI development and fostering inclusivity in biometric system evaluation.

### v3

@@
 […]
-I INTRODUCTION
-Deep learning models are primarily designed to learn specific target attributes during training; however, they frequently encode unintended, image-related attributes that can introduce significant biases into the system.
+1 Introduction
+Deep learning models are trained to learn specific target attributes, but often encode unintended image-related attributes that can adversely affect model performance and fairness.
 In the domain of biometrics, particularly face recognition, Hill et al.
@@
 [12] introduced the concept of expressivity, for face recognition, a metric that quantifies the relationship between learned network features and specific attributes, thereby enhancing the interpretability of face recognition models.
-Building on this concept, we extend the framework of expressivity to the domain of person re-identification (ReID) with the goal of evaluating how body-related features are embedded within ReID models trained primarily for identity recognition.
-Person re-identification (ReID) is a well-established research area with a range of real-world applications, including smart city infrastructure for public safety and traffic management [2, 22] and autonomous driving systems for pedestrian detection and tracking [4, 41].
-The primary objective of ReID is to accurately match and retrieve pedestrian identities across non-overlapping camera views, varying time frames, and distinct locations, all while addressing challenges such as pose variations, appearance diversity, and environmental conditions [16, 17, 46].
-Significant progress has been achieved in improving re-identification accuracy through the development of deep learning methods, which can broadly be categorized into image-based and video-based approaches.
+Building on this concept, we extend the framework of expressivity to the domain of person ReID with the goal of evaluating how body-related features are embedded within ReID models trained primarily for recognition.
+Person ReID is a well-established research area with a range of real-world applications, including smart city infrastructure for public safety and traffic management [2, 22] and autonomous driving systems for pedestrian detection and tracking [4, 44].
+The primary objective of ReID is to accurately match and retrieve pedestrian identities across non-overlapping camera views, varying time frames, and distinct locations, while addressing challenges such as pose variations, appearance diversity, and environmental conditions [16, 17, 49].
+Significant progress has been achieved in improving ReID accuracy through the development of deep learning methods, which can broadly be categorized into image-based and video-based approaches.
 Image-based ReID methods focus on selecting the most distinctive frame and extracting fine-grained spatial features, while video-based approaches aggregate temporal information across multiple frames to produce more robust identity representations.
-Recent advancements have increasingly combined these approaches, leveraging the strengths of both image-level detail and temporal consistency to achieve state-of-the-art performance.
+Recent advancements have increasingly combined these approaches, leveraging the strengths of both image-level detail and temporal consistency to achieve state-of-the-art (SoTA) performance.
 Despite these advancements, most deep learning-based ReID systems are trained to identify individuals based on visual body features, without explicitly learning specific body-related attributes.
 These models generate identity representations derived from body cues; however, similar to face recognition systems, ReID networks often unintentionally encode additional attributes related to body characteristics.
-To address this, our work conducts a comprehensive analysis of the attributes correlated with feature embeddings generated by the state-of-the-art SemReID model.
+To address this, our work conducts a comprehensive analysis of the attributes correlated with feature embeddings generated by the three SoTA ReID models and compares their robustness.
+Recently, we have come across the work of Metz et al.[29], where they also attempt to identify what information beyond identity is stored in the feature vectors from learned body recognition models.
+While they used an empirical method by training a logistic regression model to predict gender from image embeddings, it primarily demonstrates the presence of linearly separable attribute information.
+This method relies on performance metrics from a downstream classifier and does not capture the underlying statistical dependencies among attributes and representations.
+We adopt an information-theoretic perspective by applying Mutual Information Neural Estimation (MINE) to directly quantify the dependency among attribute variables and deep body recognition features.
+This allows us to measure how much information about an attribute is encoded in the feature space, regardless of classifier performance.
+Thus by moving beyond specific prediction and directly analyzing feature–attribute dependencies, our method offers a more reliable and theoretically grounded evaluation of attribute leakage and representational bias understanding.
 As ReID systems are increasingly deployed in real-world applications, there is a growing demand for explainable and transparent models.
@@
 The following are the conceptual and experimental contributions of our paper:
-  • We present the first investigation into the encoding of body attributes within the layers of a large-scale Vision Transformer (ViT)-based foundation model for person ReID.
-To enhance the interpretability of large-scale deployable ReID systems, we propose a novel post-hoc framework that explains how internal representations influence identity predictions.
-This achievement underscores the robustness of our approach, despite the inherent complexity of the model and the diversity of the dataset.
-  • In the final attention layer of the SemReID network, we observe the following order of expressivity for body attributes: BMI $>$ Pitch $>$ Yaw $>$ Gender.
+  • We present the first investigation into the encoding of body attributes within the layers of a large-scale Vision Transformer (ViT)-based self-supervised foundation model called semReID[20].
+We also compare our results against two other transformer based benchmark models.
+To enhance the interpretability of large-scale deployable ReID systems, we present a post-hoc framework that explains how internal representations influence identity predictions.
+This underscores the robustness of our method, despite the inherent complexity of the model and the diversity of the dataset.
+  • In the final attention layer of the SemReID network as well as the other networks, we observe the following order of expressivity for body attributes: BMI $>$ Pitch $>$ Gender$>$ Yaw.
 This ranking highlights the varying degrees of influence that different attributes have on the network’s predictions.
@@
 This layer-wise and temporal analysis offers deeper insights into the embedding of body attributes and their impact on ReID performance.
-II RELATED WORKS
-Person re-identification (ReID) is a critical task in computer vision that focuses on matching individuals across non-overlapping camera views, often under varying conditions such as illumination, clothing, pose, and occlusion [16, 17, 21, 46].
-Numerous works have been proposed to address this challenge in various ReID domains, such as Clothes-Changing ReID (CC-ReID)[16], video ReID[5, 19, 43, 45, 42], unconstrained ReID[11, 27, 31, 30, 47], and short-term ReID[9, 45, 40, 48].
-Among these methods, the SemReID [20] approach achieves state-of-the-art (SoTA) performance across all four domains.
-Though the interpretability of these specific models have not been extensively explored in prior literature, several works have examined the explainability of recognition algorithms in general, primarily focusing on face recognition.
-The study of bias and interpretability in biometrics has long been a subject of significant interest [36, 13, 38, 32, 33].
+2 Related Works
+Person ReID aims to match individuals across non-overlapping camera views under challenging conditions such as illumination, clothing, pose, and occlusion [16, 17, 21, 49].
+Extensive efforts have addressed this problem across domains like Clothes-Changing ReID (CC-ReID) [16], video ReID [5, 19, 46, 48, 45], unconstrained ReID [11, 28, 33, 32, 50], and short-term ReID [9, 48, 42, 51].
+Among these, SemReID [20] achieves SoTA performance across all four domains.
+While ReID interpretability remains underexplored, broader recognition systems, especially face recognition have received more attention.
+Bias and interpretability in biometrics have long been studied [38, 13, 40, 34, 35].
 Schumann et al.
-[35] introduced an auxiliary network to enhance the performance of ReID systems by incorporating complementary information into CNN features.
-More recently, Myers et al.
-[28] demonstrated that linguistic and non-linguistic representations of body shape can provide complementary identity information, improving identification in specific applications.
-These methods rely on the change in predictions with respect to a concept or attribute to interpret a network’s sensitivity to the attribute.
+[37] used an auxiliary network to enrich CNN features, and Myers et al.
+[30] leveraged both linguistic and non-linguistic body representations for identity prediction.
+These works analyze model sensitivity to attributes via concept-based prediction changes.
 Yin et al.
-[44] introduced a spatial activation diversity loss to maintain interpretability in face recognition networks during training.
-Similarly, Kim et al.
-[23] proposed a generative approach utilizing representative exemplars (prototypes) to enhance model interpretability.
-However, as noted in [24], these methods are limited to models trained from scratch and cannot be applied to pre-trained networks or those already deployed in practice.
-Post-hoc interpretability methods have also been explored to understand trained models, with TCAV (Testing with Concept Activation Vectors) [24] being one of the most influential approaches in this domain.
-TCAV interprets a model’s behavior by analyzing its sensitivity to user-defined concepts and achieves this by learning Concept Activation Vectors (CAVs) through training a linear classifier to distinguish activations produced by concept examples.
-While TCAV is effective for discrete physical concepts like color or texture, it is not suitable for assessing sensitivity to continuous or abstract attributes, such as pose angle or BMI.
-This limitation arises because generating negative examples (images without the concept) is challenging for omnipresent and continuous attributes.
-Furthermore, TCAV requires test images to belong to one of the training classes, as its computations rely on changes in logits for specific classes, rendering it incompatible with scenarios involving unseen subjects or faces.
-Other post-hoc methods, such as [1], utilize linear classifiers to analyze intermediate network layers, while Koh and Liang [25] proposed influence functions to measure model sensitivity to infinitesimal perturbations in training data.
-However, these methods are unsuitable for evaluating sensitivity to physical attributes like pose or orientation, as they primarily focus on localized perturbations.
-Saliency- and attention-based approaches [37, 8] offer another class of interpretability methods, generating attention maps to visualize spatial regions influencing model predictions.
-For person ReID, Chen et al.
-[10] introduced a learnable, pluggable interpreter for CNN-based models, which decomposes image-pair distances into attribute-based contributions and visualizes attention maps for discriminative attributes.
-While effective, this method relies on attribute-guided metric distillation, involves computationally intensive post-hoc operations, and is not adapted for transformer-based backbones.
-Moreover, while saliency maps can highlight spatial regions, they cannot assess sensitivity to abstract or non-localized concepts, such as BMI or pitch/yaw.
-Few studies explicitly address the interpretability of face recognition models on the basis of attribute hierarchy, for example Hill et al.
-[18] demonstrated a hierarchy within facial feature representations, showing that identity is nested under sex, illumination is nested under identity, and viewpoint is nested under illumination.
-Dhar et al.
-[12] further extended this analysis by conducting an expressivity-based evaluation of facial attributes, identifying a hierarchy where age has the greatest influence, followed by sex and yaw.
-We propose the use of expressivity as a measure for person ReID systems, quantifying the predictability of an attribute within a given set of features extracted using the model.
-Unlike prior methods, expressivity can be computed for both categorical and continuous attributes, allowing for a direct comparison of attribute predictability.
-Additionally, expressivity can be calculated agnostic to the ReID network backbone.
-In this work, we demonstrate results using a SoTA Vision Transformer (ViT)-based ReID model.
-This framework provides valuable insights into how body-related features are embedded within ReID models and their influence on performance, laying the groundwork for more interpretable and explainable ReID systems.
-III Proposed Method
-Our approach attempts to find the correlations between the learnt features by a state-of-the-art (SoTA) body recognition model and sensitive attributes, as highlighted in Figures 1 and 2.
+[47] introduced a spatial activation diversity loss to preserve interpretability in face recognition, while Kim et al.
+[23] proposed a prototype-based generative model.
+However, as noted in [24], such methods are limited to models trained from scratch and do not generalize to deployed networks.
+Post-hoc interpretability methods offer alternatives, notably TCAV [24], which measures sensitivity to user-defined concepts via Concept Activation Vectors (CAVs) learned through linear classification.
+While effective for discrete attributes like color or texture, TCAV struggles with continuous or omnipresent attributes (e.g., BMI, pose), where defining negative examples is difficult.
+TCAV also requires test images to belong to seen classes, limiting use in open-set scenarios.
+Other methods include layer-wise linear probes [1], influence functions [25], and saliency-based approaches [39, 8].
+For ReID specifically, Chen et al.
+[10] proposed a pluggable interpreter that attributes image-pair distances to visual cues but depends on metric distillation and is tailored to CNNs.
+Saliency maps, while helpful for spatial focus, cannot explain abstract or non-localized attributes.
+Studies in face recognition have further examined attribute hierarchies.
+Hill et al.
+[18] revealed that identity representations are nested under sex, illumination, and viewpoint, while Dhar et al.
+[12] used expressivity-based evaluations to identify a hierarchy where age dominates, followed by sex and yaw.
+We propose expressivity as a general framework to assess person ReID systems by quantifying how well an attribute can be predicted from learned features.
+Unlike prior approaches, expressivity applies to both categorical and continuous attributes and is agnostic to model backbone.
+We demonstrate its utility using ViT-based ReID models, offering insights into how body-related features are embedded and their impact on model performance, paving the way for more interpretable and explainable ReID systems.
+Figure 1: Integrating the MINE block with the ViT based SemReID [20] backbone to compute the expressivity of features with respect to attributes such as BMI, gender, pitch and yaw.
+The internal structure of the MINE block employs a simple MLP with two hidden layers to compute the expressivity of $m$-dimensional features $F$.
+By augmenting these features with an attribute vector $A$, the input to the network is extended to $(m+1)$-dimensions.
+All subjects involved provided informed consent for their participation, including the use of their images in research publications and figures.
+3 Proposed Method
+Our method as seen in Figure 1 tries to find the correlations between the learnt features by body recognition models and attributes.
 The predictability of attributes from a given set of body descriptors reflects the amount of attribute-relevant information encoded within those descriptors.
@@
 $I(X;Z)=D_{KL}\left(\mathbb{P}_{XZ}\|\mathbb{P}_{X}\otimes\mathbb{P}_{Z}\right)$ (2)
-III-A Problem Setup
-Figure 1: Integrating the MINE block in the ViT based SemReID [20] backbone to compute the expressivity of features with respect to attributes such as BMI, gender, pitch and yaw.
-The internal structure of the MINE block is shown in the next figure.
-Figure 2: The internal structure of the MINE block employs a simple MLP with two hidden layers to compute the expressivity of $m$-dimensional features $F$.
-By augmenting these features with an attribute vector $A$, the input to the network is extended to $(m+1)$-dimensions.
+3.1 Problem Setup
+Figure 2: Attribute distribution and counts in the BRIAR dataset indicate sufficient variation across the attributes of interest.
 Our dataset comprises body images of different individuals captured under varying conditions and at different distances.
@@
 We denote the set of learned feature descriptors as $\mathbf{F}$, and the corresponding sensitive attributes as $A$.
-The primary objective of this analysis is to quantify and explore the correlations between $\mathbf{F}$ and $A$ using Mutual Information Neural Estimation (MINE).
+The primary objective of this analysis is to quantify and explore the correlations between $\mathbf{F}$ and $A$ using MINE.
 Specifically, we aim to estimate the MI, denoted as $I_{\theta}(\mathbf{F},\mathbf{A})$, to gain insights into how effectively the learned features capture attribute-relevant information.
-To achieve this, each image $x_{i}$ undergoes a series of preprocessing steps before being passed through the SemReID model.
+To achieve this, each image $x_{i}$ undergoes a series of preprocessing steps before being passed through the person ReID models.
 The model extracts feature descriptors $f_{i}$, where $f_{i}\in\mathbb{R}^{m}$, representing the encoded identity and attribute information for each image.
@@
 This analysis provides valuable insights into the relationship between the network’s internal representations and sensitive attributes, helping to understand potential biases and attribute-specific influences in the model’s learned features.
-III-B Attributes and Their Relevance
+3.2 Attributes and Their Relevance
 We compute the expressivity of four annotated attributes: $g$, $BMI$, $p$ and $y$ in the extracted features.
-In Figure 3, we verify that the dataset we utilize shows enough variation with respect to these attributes, so that we can ensure that expressivity (which is a lower bound estimate of MI) is an accurate model for the corresponding attributes.
+In Figure  2, we verify that the dataset we utilize shows enough variation with respect to these attributes, so that we can ensure that expressivity (which is a lower bound estimate of MI) is an accurate model for the corresponding attributes.
 When considering g, the vector $\mathbf{A}$ is a discrete vector having a value of 1 if the gender is male and 0 if female while for $BMI$, y and p (in degrees) the values of the vector values are continuous.
-These attributes play a vital role in person ReID tasks, as they influence the expressivity of the model’s learned features.
-Attributes like $BMI$ or pose are particularly challenging to disentangle from identity for this task, making them ideal for analyzing feature relevance.
-Figure 3: Attribute distribution and counts in the BRIAR dataset indicate sufficient variation across the attributes of interest.
-III-C Expressivity of Body Features
-Understanding the expressivity of learned features in deep networks is critical for tasks that rely on nuanced feature representations, such as person re-identification (ReID).
+These attributes play a vital role in person ReID tasks, as they influence the model’s learned features.
+3.3 Expressivity of Body Features
+Understanding the expressivity of learned features in deep networks is critical for tasks that rely on nuanced feature representations, such as person ReID.
 Tishby and Zaslavsky [41] introduced the concept of utilizing MI as a quantitative metric to assess how well information is retained or transformed across the layers of a deep network.
@@
 To overcome this, Belghazi et al.
-[3] proposed Mutual Information Neural Estimation (MINE), a scalable framework that approximates MI using a neural network.
-This approach bypasses the need for explicit density computations by optimizing a neural network-based lower bound of MI, making it suitable for high-dimensional and complex datasets.
+[3] proposed MINE, a scalable framework that approximates MI using a neural network.
+This bypasses the need for explicit density computations by optimizing a neural network-based lower bound of MI, making it suitable for high-dimensional and complex datasets.
 The MI between learned feature descriptors $\mathbf{F}$ and sensitive attributes $\mathbf{A}$ is a crucial metric in evaluating the expressivity of the learned features.
-In the context of this work, $\mathbf{F}$ represents the feature embeddings produced by the SemReID model, while $\mathbf{A}$ denotes associated sensitive attributes such as gender, pose, and identity.
-The MI is mathematically defined as:
-$I_{\theta}(\mathbf{F},\mathbf{A})=\sup_{\theta\in\Theta}\mathbb{E}_{P_{FA}}\left[T_{\theta}(f,a)\right]-\log\mathbb{E}_{P_{F}\otimes P_{A}}\left[e^{T_{\theta}(f)}\right],$ (3)
-where $T_{\theta}(f,a)$ is a neural network parameterized by $\theta$, designed to approximate the MI.
+In the context of this work, $\mathbf{F}$ represents the feature embeddings produced by the ReID model, while $\mathbf{A}$ denotes associated sensitive attributes such as gender, pose, and identity.
+The MI approximate is mathematically defined as $I_{\theta}(\mathbf{F},\mathbf{A})=\sup_{\theta\in\Theta}\mathbb{E}_{P_{FA}}\left[T_{\theta}(f,a)\right]-\log\mathbb{E}_{P_{F}\otimes P_{A}}\left[e^{T_{\theta}(f)}\right]$, where $T_{\theta}(f,a)$ is a neural network parameterized by $\theta$, designed to approximate the MI.
 The joint expectation $\mathbb{E}_{P_{FA}}\left[T_{\theta}(f,a)\right]$ measures the network’s output when conditioned on the true joint distribution of features and attributes.
 In contrast, the term $\log\mathbb{E}_{P_{F}\otimes P_{A}}\left[e^{T_{\theta}(f)}\right]$ normalizes the MI estimate to ensure it captures only the dependency between $\mathbf{F}$ and $\mathbf{A}$, excluding any bias from their marginal distributions.
-Computational Steps for Mutual Information Estimation
+Computational Steps for MI Estimation:
 Step 1: Joint Expectation Approximation.
@@
 Step 3: Objective Function Formulation.
-The MI lower bound is approximated as the difference between the joint and marginal expectations, yielding the following objective function:
+The MI lower bound is approximated as the difference between the joint and marginal expectations, yielding the objective function:
 $V(\theta)=\frac{1}{b}\sum_{i=1}^{b}T_{\theta}(f_{i},a_{i})-\log\left(\frac{1}{b}\sum_{i=1}^{b}e^{T_{\theta}(f_{i})}\right).$ (5)
@@
 By iteratively computing joint and marginal expectations and updating $\theta$, the MI provides a robust metric to quantify how much attribute-relevant information is encoded in the features.
-At convergence, the MI value reflects the extent to which the model’s representations capture sensitive attribute information, offering insights into the expressivity and fairness of the learned features.
+At convergence, it reflects the extent to which the model’s representations capture sensitive attribute information, offering insights into the expressivity and fairness of the learned features.
+4 Experiments
+4.1 Dataset and Settings
+Figure 3: Attribute annotated exemplar images from the BRIAR dataset.
+All subjects involved provided informed consent for their participation, including the use of their images in research publications and figures.
+We use the BRIAR 1–5 dataset [11], a large-scale unconstrained person re-identification benchmark comprising over 1 million images and 40,000 videos captured under real-world conditions, including varying clothing, distances (100m–1km), altitudes (e.g., UAV), and environmental challenges like occlusion, blur, and turbulence.
+BRIAR includes five progressively complex subsets (BRIAR-1 to 5), increasing in identities, distractors, and capture variability.
+For our study, we extract 704,999 frames from 382,229 images and 170,522 videos, covering 2,077 unique identities (887 male and 1,190 female subjects).
+Figure 3 shows examples of images and attribute annotations from our curated subset.
+4.2 Integration of MINE with ReID Models
 Algorithm 1 Expressivity Computation on learnt representations
@@
 11:return $\text{Expressivity}\leftarrow\text{Average}(E)$
-IV EXPERIMENTS
-IV-A Dataset and Settings
-Figure 4: Attribute annotated exemplar images from the BRIAR dataset.
-We use the BRIAR 1-5 dataset [11] which has 382,229 images and 170,522 videos of which we extract 704,999 frames of 2,077 unique identities for our work.
-In our curated data subset, there are 887 males and 1,190 unique female subjects.
-BRIAR is a massive unconstrained person re-identification dataset containing over 40,000 videos and 1 million images captured across challenging real-world conditions, including varying clothes, distances (100m, 200m, 400m, 500m, 600m, 800m, 1km), altitudes (close range to UAV), and environmental factors like occlusion, blur, and physical turbulence.
-The dataset consists of 5 variants (BRIAR-1, 2, 3, 4, and 5), each with increasing complexity in terms of number of identities, distractors, and capturing conditions, making it a comprehensive benchmark for testing re-identification systems in unconstrained scenarios.
-Some example images from the dataset along with the required attribute annotation is shown in Figure 4.
-IV-B Integration of MINE with SemReID
-SemReID is a self-supervised person re-identification model that introduces a novel Local Semantic Extraction (LSE) module.
-This module leverages keypoint predictions to guide an interactive segmentation model, Segment Anything Model (SAM), in extracting precise local semantic masks for various body parts.
-These masks enable the model to isolate and extract fine-grained biometric features that enhance its ability to discern unique identities.
-Trained using a teacher-student framework, the SemReID model incorporates multiple loss functions to ensure the learned representations are robust and transferable.
-This approach simplifies the inference process by utilizing only the teacher network, paired with a single linear layer, to perform efficient re-identification.
-During inference, the model processes input data—including both still images and video frames—through the teacher encoder.
-By concatenating global and local features, the SemReID model effectively generalizes across diverse re-identification domains, negating the need for domain-specific adaptations or fine-tuning.
-To integrate MINE into the SemReID pipeline, an auxiliary neural network is employed to estimate MI by maximizing the Donsker-Varadhan (DV) lower bound.
-As described earlier, MINE provides a robust framework for estimating MI between learned feature representations and sensitive attributes, such as gender, pose, or identity.
-The implementation of this integration is detailed in Algorithm 1, which outlines the process for computing the expressivity of the learned representations.
-The process begins with a dataset consisting of $n$ images, each associated with a corresponding sensitive attribute vector $\mathbf{A}\in\mathbb{R}^{n\times 1}$.
-Features for these images are extracted from a specific layer $L$ of the SemReID model after a training epoch.
-These features, denoted as $\mathbf{F}=[f_{1},f_{2},\dots,f_{n}]^{T}$, represent both global and local information captured by the model.
-To prepare the input for the MINE network, the features and attributes are concatenated to form an augmented matrix $\mathbf{X}=[\mathbf{F}|\mathbf{A}]$.
-The MINE network $T_{\theta}$ is then initialized based on the dimensions of $X$.
-The network architecture, described in Figure 2, is a multi-layer perceptron (MLP) comprising two hidden layers with 256 and 128 units, respectively.
-These layers use Exponential Linear Unit (ELU) activations, chosen for their ability to ensure stable training and mitigate issues such as vanishing gradients.
-The MLP computes $T_{\theta}$, which is used to estimate the MI by maximizing the DV lower bound.
-To compute the expressivity of the learned representations, the process iterates for $M$ iterations.
-In each iteration, the MINE network processes the augmented matrix $\mathbf{X}$ to compute an expressivity score $e$.
-This score reflects the MI between the feature representations and the sensitive attributes.
-The scores from all iterations are stored in a list $E$, and the final expressivity measure is obtained by averaging these scores, as described in Algorithm 1.
-This integration of MINE with SemReID provides a principled approach to evaluating the expressivity of learned representations.
-By quantifying the MI between features and sensitive attributes, it becomes possible to assess the extent to which the model captures attribute-relevant information.
-Furthermore, the iterative nature of the process ensures robust estimates of expressivity while minimizing biases introduced by stochastic factors during training.
-Importantly, this integration is computationally efficient, as the MINE network operates on features extracted from the SemReID model, making it scalable.
-IV-C Hierarchical and Temporal Analysis of Attribute Influence
-To achieve a comprehensive end-to-end analysis of attribute influence within our framework, we evaluate feature and attribute correlations both hierarchically across the model’s layers and temporally throughout the training process.
-The SemReID model employs a Vision Transformer (ViT) backbone with 12 attention layers, offering a powerful architecture for capturing both global and local semantic information.
-For the hierarchical analysis, we extract features from multiple attention layers, specifically layers 2, 4, 6, 9, and 12.
-These layers are chosen to provide a granular understanding of how attribute correlations evolve at different depths of the network.
-By analyzing the intermediate layers, we gain insights into the progression of feature learning, starting from the lower layers that capture basic patterns and spatial information to the deeper layers where high-level semantic representations emerge.
-This hierarchical evaluation allows us to observe the model’s ability to encode sensitive attributes at various stages of the feature extraction pipeline and how these representations are refined.
-For the temporal analysis, we assess the evolution of feature-attribute correlations across training epochs.
-Building on prior work that established 11 epochs as the optimal training duration for identity recognition, we extend this analysis to include intermediate epochs.
-Specifically, we extract features from layers 2, 4, 6, 9, and 12 at training epochs 1, 3, 5, 8, and 11.
-This temporal sampling strategy allows us to investigate the dynamic changes in feature-attribute correlations as the model progresses through training.
-By observing earlier epochs, such as epochs 1 and 3, we capture the initial stages of representation learning, where the model begins to encode attribute information.
-In contrast, later epochs, such as 8 and 11, reveal how the model converges and solidifies its understanding of attributes in its learned representations.
-This dual analysis framework examining both the hierarchical layers and temporal progression offers a detailed perspective on how attribute information is processed and encoded within the SemReID model.
-Hierarchical analysis highlights the architectural contribution of different layers to feature representation, while temporal analysis sheds light on the dynamic learning trajectory of attribute representations over time.
-Together, these analyses enable us to uncover trends and patterns in attribute encoding, providing valuable insights into the model’s learning process in mitigating or amplifying attribute correlations.
-IV-D Implementation Details
-We initialize the MINE approximation network based on the input dimensions of the augmented matrix.
-As outlined earlier, the network is trained to estimate the lower bound of the MI between features $F$ and attribute $A$.
-For this purpose, we employ a simple multi-layer perceptron (MLP) architecture, illustrated in Figure 2, to compute $T_{\theta}$.
-The MLP consists of two hidden layers with 256 and 128 units, respectively, followed by ELU activation functions.
-This architecture is used consistently across all experiments.
-The MINE network is trained using the Adam optimizer with a learning rate of 0.001 and a batch size of 100.
-When different sets of features are used for $F$, only the input layer’s dimensions are adjusted to match the feature dimensionality.
-The network is trained until the loss in Equation (7) converges, and the expressivity is calculated as described in Algorithm 1.
-For all experiments, we set $M=16$.
-SemReID uses Vision Transformer (ViT) [14] variants as the backbone.
-Images are processed at $384\times 128$ resolution with a single forward pass.
-The inference pipeline uses a dual-stream semantic parsing strategy, where global features are processed at $768$ dimension and local semantic features at $3\times 768$ dimension where $3$ is the number of local areas.
-In our features they are face, upper body, and lower body.
-The final local features is the average of all local areas.
-We implement multi-crop feature augmentation [6, 7] using a cascade of $M=2$ global views and $N=3$ local views, followed by $L=12$ cross-attention layers for feature alignment.
-The identity embedding is computed through a BN layer, optimized for minimal latency while maintaining discriminative capabilities.
-We extract the features from this trained network which have a dimension of 1536 and pass it through the MINE block after concatenating it with the attribute vector to estimate MI.
-V RESULTS AND DISCUSSIONS
-We present our observations on the correlation between features and attributes in the SemReID model using MINE.
-The discussion is structured into three subsections: the first examines feature-attribute correlations within the hierarchical feedforward pass, while the second analyzes the evolution of these correlations throughout the training process and the third explains the advantages of our method.
-TABLE I: Summary of values for Gender, BMI, Yaw, and Pitch across different layers.
-Layer Gender BMI Yaw Pitch
-Attention Layer 2 $6.57\times 10^{-3}$ 0.005 0.13 0.42
-Attention Layer 4 0.003 0.149 0.283 0.523
-Attention Layer 6 0.004 0.72 0.73 0.73
-Attention Layer 9 0.005 0.77 0.73 0.73
-Attention Layer 12 0.03 1 0.10 0.41
-Figure 5: Expressivity trend of gender, yaw, pitch and BMI in input image over layer-wise learnt features from SemReID.
-V-A Layer-Wise Progression in Feedforward Processing
-It is crucial to emphasize that the model was trained exclusively on identity labels, without explicit supervision for BMI, gender, or pose attributes.
-Given that our architecture employs a ViTbase backbone, it consists of 12 transformer encoder blocks, each containing an attention layer.
-To analyze the evolution of feature representations, we compute the MINE at various depths—specifically after layers 2, 4, 6, 9, and 12—as summarized in Table 1 and depicted in Figure 5.
-The key observations are as follows:
-  • In the initial layers (e.g., Layer 2), MI between the learned representations and pitch/yaw attributes is relatively high, whereas gender and BMI exhibit negligible expressivity.
-This suggests that lower layers predominantly encode coarse spatial and geometric features, which are crucial for pose estimation.
-  • As we progress deeper into the ViT_base network, BMI expressivity increases rapidly, suggesting that this is a fine grained feature and is important for the model.
-By Layer 6, all attributes exhibit substantial MI with the learned features.
-However, gender expressivity remains comparatively lower, implying that it is not significantly encoded in the model’s learned representations.
+We analyze attribute correlation trends over three SoTA person ReID models namely Pose-guided Feature Disentangling (PFD)[43], Dc-former[27] and SemReID [20].
+PFD introduces a transformer-based framework for occluded person Re-ID, using pose keypoints to guide feature disentanglement between visible and occluded body regions.
+It leverages a ViT-base backbone to encode patch-level features and a transformer decoder with learnable semantic tokens aligned to pose-guided regions via a Pose-View Matching module.
+A pose-guided push loss further encourages discrimination of visible parts while suppressing noisy occluded features for robust identity matching.
+The DC-Former utilizes a ViT-small backbone to improve person ReID by learning multiple diverse and compact embedding subspaces.
+It uses multiple class tokens within the Vision Transformer framework to represent distinct embedding spaces.
+A Self-Diverse Constraint (SDC) is applied to encourage these embeddings to be both diverse and compact, improving the model’s ability to distinguish between similar identities.
+Additionally, a Dynamic Weight Controller (DWC) balances the importance of each embedding space during training, leading to robust and discriminative feature representations.
+SemReID is a self-supervised person ReID model that introduces a novel Local Semantic Extraction (LSE) module.
+This module uses keypoint predictions to guide the Segment Anything Model (SAM), producing precise local semantic masks for various body parts.
+These masks allow for the extraction of fine-grained biometric features, enhancing identity discrimination.
+SemReID is trained using a teacher-student framework with multiple loss functions to promote robustness and transferability.
+At inference, only the teacher network and a single linear layer are used, enabling efficient re-identification by processing input data through the teacher encoder.
+Global and local features are concatenated to facilitate generalization across domains without domain-specific fine-tuning.
+To evaluate feature expressivity, MINE is integrated into these pipelines as an auxiliary neural network estimating mutual information (MI) by maximizing the Donsker-Varadhan (DV) lower bound.
+Given a dataset of $n$ images with corresponding sensitive attributes $\mathbf{A}\in\mathbb{R}^{n\times 1}$, features are extracted from a specific layer $L$, yielding $\mathbf{F}=[f_{1},f_{2},\dots,f_{n}]^{T}$.
+These features, capturing both global and local cues, are concatenated with attributes to form the input matrix $\mathbf{X}=[\mathbf{F}\,|\,\mathbf{A}]$.
+The MINE network $T_{\theta}$, a multi-layer perceptron (MLP) with hidden layers of 512 and 128 units and ELU activations, is initialized based on $\mathbf{X}$’s dimensions.
+Over $M$ iterations, $T_{\theta}$ processes $\mathbf{X}$ to compute expressivity scores $e$, which estimate the MI between features and attributes.
+These scores are stored in a list $E$, and the final expressivity is computed as their average (Algorithm 1).
+This integration provides a principled and scalable approach to quantify attribute-relevant information in learned representations.
+The iterative MINE process ensures stable and unbiased estimates of expressivity while remaining computationally efficient, as it operates on pre-extracted features from these models.
+4.3 Hierarchical and Temporal Analysis of Attribute Influence
+To comprehensively analyze attribute influence in our framework, we examine feature–attribute correlations both hierarchically across model layers and temporally over training epochs.
+The SemReID model and the PFD model use a ViT backbone with 12 attention layers, while the Dc-former model uses 8 attention layers to capture rich global and local semantics.
+Hierarchical Analysis: We extract features from layers 2, 4, 6, 9, and 12 for the PFD and SemReID models while features from layers 2,3,4,6 and 8 for the Dc-former to study how attribute correlations evolve with network depth.
+We consider layers 2,4 to be early 6,9 to be mid and 12 to be late for ViT base while 2,3 to be early, 4,6 to be mid and 8 to be late for ViT small.
+These layers are selected to provide a fine-grained view of the learning process from early layers capturing basic spatial patterns to deeper layers encoding high-level semantics.
+Temporal Analysis: We train each of these models for 11 epochs and to assess how these correlations change during training, we analyze the final layer features at epochs 1, 3, 5, 8, and 11.
+Early epochs (1, 3) highlight the emergence of attribute encoding, while later epochs (8, 11) illustrate how these representations stabilize as the model converges.
+Together, this dual analysis, provides a detailed understanding of how attribute information is processed, encoded, and evolved within these models.
+It uncovers key trends in the model’s capacity to learn, refine, or suppress sensitive attribute correlations throughout training.
+Models Rank 1 5 10 TAR @1% @10%
+DC-Former 27.98 54.06 62.28 39.84 84.51
+PFD 32.92 55.65 75.73 47.97 71.22
+SemReID 34.84 56.95 66.33 54.09 89.03
+Table 1: Quantitative comparison between DC-Former, PFD, and SemReID for identification task (Best in bold).
+4.4 Implementation Details
+We initialize the MINE network based on the input dimensions of the augmented matrix, using a two-layer MLP (512 and 128 units, ELU activations) as seen in Fig.
+2 to compute $T_{\theta}$.
+The weights are initialized using Xavier normal initialization.
+This setup, consistent across experiments, is trained with Adam (learning rate=0.001, batch size=100) until Equation (7) converges.
+Only the input layer adapts to the feature dimensionality.
+Expressivity is calculated per Algorithm 1, with $M=5$.
+For SemReID, we use ViT variants [14] with $384\times 128$ inputs in a single forward pass.
+A dual-stream setup extracts 768-dim global and $3\times 768$-dim local semantic features (face, upper, lower body), averaged for the final local embedding.
+Multi-crop augmentation [6, 7] uses $M=2$ global and $N=3$ local views, followed by $L=12$ cross-attention layers.
+Identity embeddings are computed via a BN layer for efficiency.
+Final 1536-dim features are concatenated with attribute vectors and fed to MINE to estimate MI.
+DC-Former [27] uses multiple class tokens with SDC ($\lambda=1$) to create diverse embedding subspaces, using dynamic weight controller for token balancing.
+PFD [43] utilizes overlapping patch embedding (step size=12), pose-guided feature aggregation via HRNet-estimated keypoints (threshold $\gamma=0.2$), and part view based decoder with $N_{v}=17$ learnable semantic views; decoder layers set to 6.
+Gender Yaw BMI Pitch
+No. DCFormer SemReID PFD DCFormer SemReID PFD DCFormer SemReID PFD DCFormer SemReID PFD
+Layer early (2-s,2-b) 0.3252 0.0978 0.3250 0.3109 0.0026 0.5443 0.3588 0.0950 0.5153 0.4430 0.1461 0.6151
+early (3-s, 4-b) 0.3330 0.0999 0.4176 0.3595 0.0210 0.7181 0.4102 0.2174 0.6509 0.4733 0.2243 0.7259
+mid (4-s, 6-b) 0.3490 0.1829 0.4321 0.3458 0.0504 0.7611 0.5023 0.2816 0.7940 0.4203 0.2998 0.7678
+mid (6-s, 9-b) 0.3101 0.2150 0.4428 0.1909 0.0716 0.5369 0.6101 0.4714 1.2750 0.3440 0.3190 0.5925
+late (8-s,12-b) 0.2856 0.1586 0.4280 0.1456 0.0119 0.4074 0.6832 0.5740 1.2316 0.3347 0.2727 0.4950
+Epoch 1 0.3272 0.2115 0.4864 0.2454 0.2519 0.7119 0.4968 0.4283 0.9898 0.3065 0.3722 0.7084
+3 0.2980 0.1388 0.4403 0.1772 0.000094 0.6426 0.6198 0.3724 1.0660 0.3332 0.2419 0.5998
+5 0.2945 0.1224 0.4425 0.1756 0.000029 0.5512 0.6580 0.4194 1.1495 0.3172 0.2052 0.6097
+8 0.2914 0.1461 0.4211 0.1896 0.00145 0.4328 0.7047 0.5479 1.2280 0.3342 0.2517 0.5251
+11 0.2856 0.1586 0.4280 0.1456 0.0119 0.4074 0.6832 0.5740 1.2316 0.3347 0.2727 0.4950
+Table 2: Expressivity scores for Gender, Yaw, BMI, and Pitch across layers and training epochs for DCFormer, SemReID, and PFD.
+x-s and y-b denote the $x^{th}$ and $y^{th}$ layer of ViT small and base respectively.
+5 Results and Discussions
+On the person ReID task for BRIAR, SemReID demonstrates the strongest performance (Table 1).
+In this section, we analyze feature-attribute correlations using MINE.
+The discussion is structured into three subsections: (1) spatial trends across the hierarchical feedforward pipeline, (2) temporal progression during training, and finally (3) the advantages of our method.
+5.1 Hierarchical Feedforward Progression of Attribute Expressivity
+Figure 4: Expressivity trends of gender, yaw, pitch and BMI in input image over layer-wise learnt features from SemReID.
+To study the evolution of attribute correlations across network depth, we extract features from layers 2, 4, 6, 9, and 12 for the PFD and SemReID models, and from layers 2, 3, 4, 6, and 8 for DCFormer.
+For ViT-Base (PFD and SemReID), we categorize layers 2 and 4 as early, 6 and 9 as mid-level, and 12 as late.
+For ViT-Small (DCFormer), layers 2 and 3 are considered early, 4 and 6 as mid-level, and 8 as late.
+Figure 4 visualizes attribute expressivity per layer, and Table 2 reports corresponding values.
+Our analysis reveals the following trends:
+  • BMI Encodes Deeply Across All Models: BMI expressivity increases consistently with depth in all models.
+PFD shows the sharpest rise from around 0.6 in early layers to over 1.2 in the final block—highlighting that BMI becomes increasingly entangled in deeper feature hierarchies.
+This suggests that morphometric cues are emphasized more in later layers.
+  • Yaw Peaks in Mid-Layers and Then Drops: Yaw exhibits a peak in the mid-layers and declines in the final block across all models.
+This is most prominent in PFD, where yaw rises to $\sim 0.75$ mid-network and then drops to $\sim 0.4$.
+The pattern suggests intermediate layers encode pose cues strongly, which are later partially suppressed as identity features dominate.
+  • Pitch Follows a Similar Pattern to Yaw, But Less Drastic: Pitch also peaks at the mid-layer stage before dropping in the late layers.
+However, unlike yaw, pitch maintains moderate expressivity even in deeper layers, particularly in DCFormer and PFD.
+This highlights partial retention of pitch as a latent factor.
+  • Gender Expressivity is Moderate and Stable: Gender shows moderate expressivity across all layers, particularly in PFD and DCFormer, where its scores remain relatively consistent.
+While it does not peak as strongly as BMI or pose attributes, it is clearly encoded and not suppressed.
+In SemReID, gender shows a slight rise in mid-layers before declining in the final block.
+Overall, gender expressivity is less depth-sensitive but remains a persistent latent factor across the network hierarchy.
+  • Model-wise Trends: PFD encodes all attributes more strongly across all layers, indicating greater attribute leakage and less disentanglement.
+DCFormer exhibits moderate levels of expressivity and clear hierarchical trends.
+SemReID consistently suppresses attributes more effectively, especially yaw and gender, suggesting more robust identity representation.
 […]
-Thus, for the semReID model trained purely on identity labels, BMI emerges as the most correlated attribute, followed by pose, with gender being the least correlated.
-This conclusion aligns with intuitive human perception, as body recognition inherently relies on a person’s shape and pose as primary cues.
-From a perceptual standpoint, humans often associate body identity with physical attributes such as body shape, proportions, and posture, making BMI and pose naturally dominant in recognition tasks.
-We want to clarify at this stage that, the data processing inequality (DPI) [23] states that for three random variables $P$, $Q$, and $R$ forming a Markov chain $P\rightarrow Q\rightarrow R$, the MI satisfies $MI(P,Q)\geq MI(P,R)$.
-This principle formalizes the idea that no processing of data can increase MI.
-In the context of neural networks, let $P$ represent a random variable (e.g., BMI, pose, or identity), and $Q$ and $R$ represent features extracted at different layers of the network, where $R$ is a deterministic function of $Q$ (i.e., $R$ corresponds to a deeper layer).
-Since $R$ is derived from $Q$, $P$, $Q$, and $R$ form a Markov chain, and the DPI implies that the MI between $P$ and the features cannot increase as we go deeper into the network.
-However, the expressivity results are not strictly monotonically decreasing, which might seem to contradict the DPI.
-This apparent inconsistency arises because, in our work, expressivity refers to the alignment between feature representations and a given attribute, rather than MI in the strict sense of information theory.
-As noted in [14], the feature representations in this context are more closely tied to the predictability of an attribute rather than its information-theoretic content.
-Therefore, expressivity in this work reflects the strength of the relationship between features and attributes, rather than their theoretical MI.
-This distinction allows us to interpret the expressivity results as a measure of the model’s ability to capture attribute-specific representations, which can vary across layers.
-Figure 6: Expressivity trend of gender, yaw, pitch and BMI in input image over epoch-wise learnt features from SemReID.
-TABLE II: Summary of values for Gender, BMI, Yaw, and Pitch across different epochs.
-Epoch Gender BMI Yaw Pitch
-Epoch 1 0.012 0.87 1.3 0.84
-Epoch 3 0.402 0.305 0.009 0.004
-Epoch 5 0.004 0.76 $6\times 10^{-3}$ $4\times 10^{-3}$
-Epoch 8 0.012 0.80 0.05 0.22
-Epoch 11 0.03 1 0.10 0.41
-V-B Progression of Training
-To understand how attribute-feature correlations evolve over training, we analyze the expressivity of gender, yaw, pitch, and BMI in the final attention layer across different training epochs.
-This is summarized in Table II and visualized in Figure 6.
-At each epoch, the feature representations are 1536-dimensional, and attribute-wise expressivity is computed using Algorithm 1.
-We observe that:
+5.2 Temporal Progression of Attribute Expressivity
+Figure 5: Expressivity trends of gender, yaw, pitch and BMI in input image over epoch-wise learnt features from SemReID.
+We analyze the temporal evolution of MINE expressivity scores across epochs for the same four attributes measured at the final attention layer of the three models: SemReID, DCFormer, and PFD.
+Figure 5 shows the epoch-wise trends, and Table 2 reports the absolute scores.
+We observe the following key patterns:
 […]
-  • A significant reduction in yaw expressivity is observed, while BMI expressivity remains relatively stable.
-This suggests that yaw-related variations are gradually suppressed as the model prioritizes more identity-relevant features.
-Notably, both yaw and gender exhibit minimal correlations with the learned representations, implying that the model is learning to be invariant to these attributes in its deeper layers.
-Conversely, BMI remains a highly correlated feature, followed by pitch, reinforcing its role as a more identity-relevant attribute.
-[…]
-These findings indicate that the training process makes feature representations increasingly invariant to certain body attributes, such as yaw and gender, while preserving information related to BMI and to some extent pitch.
-V-C Advantages of Expressivity for Person ReID
-In this subsection we justify the usage of MINE over other existing methods. The key reasons can be summarized as:
+  • Yaw Suppression Over Time: All models exhibit a notable decline in yaw expressivity with training, dropping from over 0.7 to around 0.4 by epoch 11 in PFD, and collapsing to near zero after epoch 2 in SemReID.
+This indicates active suppression of yaw as models optimize for pose invariant identity representation.
+  • BMI Persistence Across Models: Unlike other attributes, BMI expressivity shows a rising or stable trend for all models.
+In PFD, it increases from 1.0 to 1.2, while DCFormer and SemReID show comparatively modest gains.
+This highlights BMI as a persistent entangled attribute, potentially acting as a confounding factor in identity representations.
+  • Gender Expressivity is Low and Stable: Gender shows the lowest expressivity throughout training across models except for yaw.
+In SemReID, it drops early and stabilizes near 0.13, indicating progressive decoupling over time.
+  • Pitch Shows Intermediate Trends: Pitch expressivity stabilizes after epoch 3 for all models, with DCFormer consistently encoding more pitch information than SemReID.
+The stability suggests that pitch is preserved during training but is less dominant than BMI.
+  • Model Comparisons: PFD consistently encodes more attribute information than the other two models across all attributes and epochs.
+In contrast, SemReID exhibits the lowest attribute expressivity overall, indicating more robust identity feature learning with reduced sensitivity to spurious cues like yaw and gender.
+These findings reveal that during training, pose-related cues like yaw are actively suppressed, particularly in SemReID.
+However, body-morphometric attributes such as BMI persist in the feature space, suggesting a need for targeted disentanglement strategies for such persistent confounders.
+5.3 Advantages of Expressivity for Person ReID
+In this subsection we further justify the usage of MINE over other existing methods. The key reasons are:
   1. 1.
-    Enables Attribute Comparison Across Features: In the context of person ReID, comparing the significance of various attributes within a network’s feature set is critical for understanding how identity information is organized.
-Expressivity facilitates this by normalizing the error rates of attributes to a common scale, as illustrated in Figures 5 and 6.
-This standardization enables direct comparison of attribute contributions, unlike conventional metrics influenced by varying scales.
+    Supports Both Discrete and Continuous Attributes: Expressivity is versatile and applicable to both discrete (e.g., gender) and continuous (e.g., pitch angle) attributes.
+For example, gender expressivity can be computed using a binary attribute vector $A$.
+Unlike TCAV [24], which is tailored to discrete attributes, expressivity naturally extends to continuous concepts such as pose or BMI where defining clear negative examples is difficult.
+This makes it especially valuable in ReID, where continuous attributes are often crucial.
   2. 2.
-    Supports Both Discrete and Continuous Attributes: This work focuses on body attributes for ReID, expressivity is versatile and can be applied to both discrete (e.g., gender) and continuous (e.g., pitch angle) attributes.
-For instance, the expressivity of a concept such as gender can be computed if a binary attribute vector $A$ (indicating 1 for male and 0 for female or vice versa) is available.
-Unlike TCAV [24], which is well-suited for discrete attributes, expressivity extends to continuous concepts like pose angle or BMI, where identifying negative examples (images where the concept is absent) is challenging.
-This capability is particularly valuable for person ReID, where continuous attributes often play a critical role.
-  3. 3.
     Independent of Training Identities: Previous methods require computing changes in logits, which limits their applicability to images belonging to training identities.
@@
 This independence makes it an effective tool for analyzing unseen attributes not explicitly included during training.
-VI Conclusion
+6 Conclusion
 We propose a method to quantify the information a ViT-based person ReID network learns about various attributes without being explicitly trained on them, by analyzing their expressivity on learnt features.
@@
 (3) Temporally, BMI expressivity increases throughout training, while yaw and pitch decline sharply, with yaw showing the steepest drop.
-Gender, notably, has minimal correlation with learned features.
-These findings highlight BMI as the most significant attribute, followed by yaw and pitch and gender for the person ReID task.
+Gender is entangled in the features moderately and has minimal variation with the evolution of learned features.
+These findings highlight BMI as the most significant attribute, followed by yaw gender and pitch for the person ReID task.
 However, since expressivity approximates MI, it is influenced by entropy and attribute label distribution, potentially affecting cross-attribute comparisons which is an inherent limitation of all MI-based approaches.
-VII ACKNOWLEDGMENTS
+7 Acknowledgements
 SH and RC are supported by the BRIAR project.
@@
 Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
-VIII Ethical Impact Statement
-This research utilizes the BRIAR dataset, which was collected by the Test and Evaluation team under strict consent guidelines and with approval from the appropriate ethical review board.
-All subjects involved provided informed consent for their participation, including the use of their images in research publications and figures.
-Our study aims to analyze how demographic and physiological covariates such as body mass index and gender may introduce biases in body-based recognition systems.
-By systematically examining these potential biases, we seek to contribute to the broader discourse on fairness, accountability, and transparency in biometric technologies.
-Understanding and mitigating bias in recognition systems is critical for ensuring equitable outcomes across diverse populations.
-This research aligns with the outlined ethical principles by advocating for responsible AI development and fostering inclusivity in biometric system evaluation.